OpenStudy (anonymous):

What degree Fahrenheit is it if the temperature is 15 degrees Celsius? Show all steps or reasoning. Be sure to check your answer for reasonableness.

OpenStudy (calculusxy):

Here's an easy method that should help you do all the calculations to convert Celsius to Fahrenheit. 1. Take the temperature and divide it by 1.8. 2. Add 32 to the quotient. The answer will then be telling you the degrees in Fahrenheit. Let's put this to work. 15/1.8=8.33(which we can round off to 9) 9+32=41 So 15 degrees Celsius equals to 41 degrees Fahrenheit. Now if you wish to convert Fahrenheit to Celsius then just do the complete opposite of why we just did now.

OpenStudy (tkhunny):

Divide by 1.8 is a non-trivial exercise. It's also incorrect. It should be MULTIPLY by 1.8. It's still a non-trivial exercise. Notice, however, that 1.8 = 9/5. Now THAT we might be able to do in our heads. 15(9/5) + 32 15/5 = 3 3*9 = 27 27 + 32 = 59 Always check standard values to ensure correctness. 40ºC = 104ºF -40ºC = -40ºF 40(9/5) + 32 40/5 = 8 8*9 = 72 72 + 32 = 104 -- Perfect